An atom has a diameter of 1.50 Å and the nucleus of that atom has a diameter of 7.50×10−5 Å . Determine the fraction of the volume of the atom that is taken up by the nucleus. Assume the atom and the nucleus are a sphere
1
Expert's answer
2020-09-08T05:58:25-0400

QUESTION # 132030

An atom has a diameter of 1.50 Å and the nucleus of that atom has a diameter of 7.50×10−5 Å. Determine the fraction of the volume of the atom that is taken up by the nucleus. Assume the atom and the nucleus are a sphere

Answers

Volume of nucleus, V=4/3∏r3

                                                    = 4/3 × 22/7 × (3.75 ×10-5)^3

                                        =2.209810-13

Volume of atom, V = 4/3 × 22/7 × (7.5 × 10-1)^3

              = 1.7679

Fraction of atom taken up by nucleus = (2.209810-13) /1.7679

                                                                    = 1.25×10-13